A regular polygon of(2k +1)sides has 140 as d size of each interior angle.find k
Answer by hammy(9)

S=180*(n-2)
(2k+1)*140=180
(2k+1)/(2k-1)=9/7
18k-9=14k+7
4k=16
k=4